A trumpet creates a sound intensity level of 52.1 dB at a distance of 3.6 m. What is the sound intensity level (in dB) of 7 such trumpets at a distance of 10.9 m?
Answer #1

Intensity level of sound and distance can be related by the equation, - 82 where, L, intensity level at - distancer, 7 = - 2
Answer #2

SOLUTION :


I2 

= intensity level of 1 trumpet at a distance of  r2.

= I1 * r1^2 / r2^2    

= 52.1 * 3.6^2 / 10.9^2.

 (I1 is intensity of a trumpet at r1 (3.6 m) away = 52.1 DB, r2 = 10.9 m)

= 5.6832 DB 


So,


Intensity level of 7 trumpets 

= 7 * 5.6832

= 39.78 DB (ANSWER).

answered by: Tulsiram Garg